Solve one using elimination and one using substitution;
1)4x-y=10
-3x+y=-6
2)5x-3y=16
2x+7y=-10

Found 2 solutions by mananth, solver91311:
Answer by mananth(16949)

4 x -1 y = 10 .............1
-3 x + 1 y = -6 .............2

Add the two equations
x = 4
/ 1
x = 4
plug value of x in (1)
4 x + -1 y = 10
16 + -1 y = 10
-1 y = 10 -16
-1 y = -6
y = 6
(4,6)
-
5 x -3 y = 16 .............1
7 y = -10 -2 x ---------2
/ 7 =
y = -1.43 -0.29 x
Substitute y in (1)
5 x -3 ( -1.43 -0.29 x ) =
5 x + 4.29 + 0.86 x = 16
5 x + 0.86 x = 16 -4.29
5.86 x = 11.71
/ 5.86
x= 2
Plug the value of x in (1)
5 x -3 y = 16
5 * 2 -3 y = 16
10 -3 y = 16
-3 y = 6
/ -3

(2,-3)



Answer by solver91311(24713)







Notice that the coefficient on in the first equation is the additive inverse of the coefficient on in the second equation. Just add the equations:



Notice that the variable (in this case) is eliminated, hence the name Elimination Method.



Substitute the calculated value back into either given equation:





The solution set is

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism